OpenStudy (anonymous):

please help? Write the slope-intercept equation for the line with slope -2/3 and y-intercept (0, 3) and graph the line

OpenStudy (anonymous):

Slope-intercept form : y=mx+b m=-2/3 b=3 y=-2/3x+3 To graph, just graph by using the y-intercept and follow the slope. Go down 2 units, then right 3.

OpenStudy (anonymous):

the formula y=mx+b works. b=3 and m= -2/3 so y=-2x/3 +3

OpenStudy (anonymous):

Its y= mx+b m=slope and b= y intercept. y= -2/3x+ 3 Thats the y-form or the slope intercept form. :)

OpenStudy (anonymous):

The equation of a line is given by \[y-y_1 = m(x-x_1)\] You have the slope (\(-\frac{2}{3})\), and a point \((0,3)\)